A spectroscope with a diffraction grating having 633.0 lines / millimeter is used to measure the angle of an emission line.

What is the spacing ( d ) between grating lines, expressed in nanometers? Report to the nearest whole number,without any units.
Answer #1

There are 633.0 lines in 1 mm of the diffraction grating.

So the spacing between slits (d)=1 mm/633.0 =0.0015797788 mm=0.0015797788 x 10-3 m

d=0.0015797788 x10​​​​6 x10-6 x 10-3 (multiply with 106 and 10-6)

d=1579.7788 x10​​​​-9 m

d1580 nm i.e. 1580 (nearest whole number without units)

(1 nm=10-9 m and 1 mm=10-3 m)